# Learning Simulatable Models of Cloth with Spatially-varying Constitutive Properties

### v2

@@
Abstract
Materials used in real clothing exhibit remarkable complexity and spatial variation due to common processes such as stitching, hemming, dyeing, printing, padding, and bonding. Simulating these materials, for instance using finite element methods, is often computationally demanding and slow. Worse, such methods can suffer from numerical artifacts called “membrane locking” that makes cloth appear artificially stiff. Here we propose a general framework, called Mass-Spring Net, for learning a simple yet efficient surrogate model that captures the effects of these complex materials using only motion observations. The cloth is discretized into a mass-spring network with unknown material parameters that are learned directly from the motion data, using a novel force-and-impulse loss function. Our approach demonstrates the ability to accurately model spatially varying material properties from a variety of data sources, and immunity to membrane locking which plagues FEM-based simulations. Compared to graph-based networks and neural ODE-based architectures, our method achieves significantly faster training times, higher reconstruction accuracy, and improved generalization to novel dynamic scenarios.
1 Introduction
@@
In this work, we introduce a learning-based framework, called Mass-Spring Net, that addresses all these problems. For run-time simulation, Mass-Spring Net uses a simple mass-spring network cloth model that is easy to implement and efficient to simulate in real-time applications. Moreover, we show that the Mass-Spring Net can effectively learn spatially-varying constitutive properties of cloth from simulated experiments using a more complex and slow model (e.g., an FEM model). Thus the Mass-Spring Net may be used as a real-time “surrogate” model for a more complex off-line model.
Even though there is a popular perception that mass-spring networks are unsophisticated and less appropriate for modeling continuum mechanics, that is not necessarily the case. As observed by Breen et al. , Hearle and colleagues at Manchester’s renowned Department of Textiles had investigated applications of shell and plate models in the 1980s, and concluded: “But, in dealing with the three-dimensional buckling of textile fabrics, neither the terminology nor the methodology of the established mechanical theory of the bending of plates and shells is of much help because of the limiting assumptions that are made” (Amirbayat and Hearle, 1989). Mass-spring networks could better capture the anisotropic behavior of cloth, and avoid numerical artifacts of FEM discretization of thin shells. We demonstrate, with a series of experiments, that the Mass-Spring Net is resistant to the membrane locking problem that plagues FEM simulations, while avoiding the complexity of special solutions. Remarkably, we observe that the Mass-Spring Net behaves better under bending than even the FEM model that it was trained with, thereby achieving superresolution in inference. We illustrate these phenomena and offer some potential explanations.
@@ 2 Related Work
Cloth Simulation and constitutive property estimation. Simulations pertaining to cloth have always been a prime subfield in graphics and simulation research due to their varied industrial applications, from yarn-level simulators by Sperl et al. (2022); Cirio et al. (2014) which specialize in woven cloths, to continuum-level or meshed-based simulators (Shao et al., 2025; Ju et al., 2024; Grigorev et al., 2023; Kairanda et al., 2023; Zhao et al., 2023; Feng et al., 2022; Li et al., 2022a; Santesteban et al., 2022; Bertiche et al., 2022; Li et al., 2022b; Pfaff et al., 2020; Liang et al., 2019; Wang et al., 2011). Many of these works, including Ju et al. (2024); Feng et al. (2022); Wang et al. (2011) focus on accurate constitutive parameter estimates by real-to-sim adaptations using specialized measurement devices; others (Shao et al., 2025; Grigorev et al., 2023; Pfaff et al., 2020) focused more on learning constitutive properties from synthetic data, and demonstrated great generalizability to real-world data (Zheng et al., 2024; Li et al., 2022a, b; Liang et al., 2019), or more specifically, learning garment deformation models jointly with human motion kinematics (Stuyck and Chen, 2023; Zhao et al., 2023; Grigorev et al., 2023; Li et al., 2024; Bertiche et al., 2022; Santesteban et al., 2022). However, cloth or garments chosen as the subject to estimate haven been almost exclusively homogeneous, composed of a single fabric without complex fine-scale patterns and structures; the only exceptions are Stuyck and Chen (2023) and Grigorev et al. (2023), yet both explores fitting specifically of garments on virtual try-on’s. The focus of our work is to address the broader problem of estimating any piece of cloth’s heterogeneous constitutive properties, such as stiffness that varies across space using synthetic data. We distinguish from work like (Stuyck and Chen, 2023; Li et al., 2022b; Pfaff et al., 2020) that our estimation pipeline does not require a reference mesh as input, and from (Zheng et al., 2024; Grigorev et al., 2023; Stuyck and Chen, 2023; Bertiche et al., 2022) that we focus on the broader domain of modelling constitutive properties of a standalone cloth, rather than the specific domain of high-quality reconstruction of garment kinematics on moving human avatars. We show that by fitting a physically correct mass-spring network using our novel loss and training curriculum, we achieve comparable or better accuracy and generalizability than previous approaches but at a significant reduction in training time and model complexity.
@@ 3 Method
We map the predicted spring forces to internal forces on each surrogate particle Liu et al. (2013),
@@
3.1 Spatial Discretization and Resampling
Given a desired number of surrogate particles , we build our surrogate mass-spring network as follows: we isometrically unroll the source system’s cloth specimen into a rectangle in the plane, then discretize this rectangle with a -vertex regular square grid, as shown in Fig. 5 (a). We place a surrogate spring at each edge of the grid, with the spring rest length determined by the edge length in this rest configuration.
As shown in Fig. 5 (b), the source system landmarks do not necessarily correspond to the surrogate particles (and the resolution of the surrogate particle grid might be much coarser or finer than the density of landmark points). Therefore we must resample the given landmark positions at each frame to positions of the surrogate particles.
@@ 3.1 Spatial Discretization and Resampling
3.2 Neural Constitutive Modeling
@@ 3.3 Simulation and Training
Many alternative loss functions for training the surrogate could also be used, such as position loss (Ma et al., 2023) or acceleration loss (Sanchez-Gonzalez et al., 2020), but in our ablation experiments (Sec. C.2) we found that the impulse loss works best in our setting.
@@
4 Experiments
4.1 Synthetic Data Generation
Constructing source systems. We build two source systems of cloths modeled under different simulation frameworks: (i) A square-shaped cloth made up of particles and springs. Forces that act upon particles come from springs only; (ii) A square-shaped cloth constructed from a triangular mesh. Internal forces are evaluated based on FEM principles. We use NVIDIA WARP’s FEM-based cloth simulation publicly available on Github (Macklin, 2022a). We discretize each cloth into a particle grid. The choice of varies across experiments; for the experiment in which we compare Mass-Spring Net with baselines on generalizability to novel boundary and initial conditions, we use , which is close to the discretization scale picked by recent work such as Shao et al. (2025). We assign different stiffness values to either springs in the mass-spring cloth or triangles in the mesh-based cloth so they have spatially-varying constitutive properties: see Fig. 1.
Generating rollouts with different initial and boundary conditions. To collect ground-truth simulation data, for each cloth we anchor a set of its vertices so they cannot move, and we simulate the deformation of the rest of the cloth over time. The set of vertices being fixed in space, formally defined as anchor neighborhood defines the boundary condition of a rollout; the initial position and velocity of all particles defines its initial condition. For each rollout , we randomly pick two anchor neighborhoods as boundary condition , and we apply a random rotation to the sheet to assign the cloth with initial condition : see Fig. 6. For training the Mass-Spring Net and the Neural ODE baseline we use rollouts; for testing we use rollouts. Each rollout lasts seconds, during which the cloth falls under the influence of gravity and Rayleigh damping force , where Rayleigh damping constant . We simulate with timestep size of , and this time discretization scheme gives us time steps of data from each rollout.
4.2 Metrics
In some of our experiments, we evaluate Mass-Spring Net by applying it to a source system that is itself a simulation of a mass-spring network. In this setting, we can easily compare the surrogate’s estimated spring stiffness and damping parameters to those of the source system. We quantify the error in the reconstructed parameters by computing the root-mean-squared error (RMSE) between their estimated and ground-truth values:
@@ 4.2 Metrics
4.3 Reconstructing Spatially Varying Materials
The goal of this experiment is to show that Mass-Spring Net can accurately estimate the stiffness and damping properties of a source cloth with spatially varying stiffness. We use the and metric established in Eq. 13 to assess the quality of our constitutive property estimates, and to use them, we build the source cloth with masses and springs, and we construct a surrogate system with the same resolution as the source cloth. Bending and the complete set of shear springs are included in both the source and the surrogate. This allows us to establish a one-to-one mapping from springs in the source cloth to springs in the surrogate cloth. We also assess the quality of motion reconstruction using the Motion RMSE metric from Eq. 14.
Fig. 7 shows the accuracy of stiffness and damping estimates, and the reconstruction error obtained from a test set which contains -long rollouts. The source cloth we used to generate training and test rollouts is a square-shaped, cloth discretized into a particle grid connected by springs. The source cloth contains springs with three stiffnesses: , and , and all springs share the same damping constant of . Fig. 7 (a) tells that most of our estimated stiffnesses are close to the ground-truth. The damping estimates are also close to the ground-truth on average, despite some variations: see (b).
4.4 Immunity to Membrane Locking
@@
To further validate the super-resolution effect of the mass-spring-based surrogate, as shown in Fig. 2 (b), (c), we pin two corners of a square-shaped FEM-simulated mesh along a diagonal, make the cloth fold under the influence of gravity and Rayleigh damping, and sample the cloth’s equilibrium configurations. We compare the equilibrium configurations under two cases: (i) “Aligned”, i.e. folding along the major diagonal that coincides with diagonal edges in the mesh; (ii) “Misaligned”, i.e. folding along the minor diagonal. Under the misaligned case, as we expect, the cloth sags less than the aligned case, and the difference in equilibrium configurations is more apparent on the low-resolution mesh. As shown in Fig. 2 (d), our mass-spring-based surrogate which includes the full set of diagonal and bending springs, and trained on landmark trajectories sampled from the low-resolution mesh in (b), however is immune to this effect, because we can see that at equilibrium configuration it is closer to the high-resolution mesh that suffers less from membrane locking. Further, by comparing Fig. 2 (d), (e) we can see that the surrogate is immune to the pseudo-stiffness caused by membrane locking not by chance, but because it can properly fit to the actual bending resistance of the mesh— the surrogate at equilibrium configuration is closer to the mesh with bending stiffness (d) than the mesh with zero bending stiffness (e).
@@
While both Mass-Spring Net and MGN are data-driven methods that encode cloth’s constitutive properties and reconstruct reference motion in inference, what makes Mass-Spring Net stands out is its robustness to numerical artifacts that exist in the training data. We believe Mass-Spring Net attains the robustness by more accurately estimating the cloth’s constitutive properties, after seeing trajectories sampled in a multitude () rollouts, many of which would not lead to a “locked” equilibrium configuration as in the misaligned case. MGN on the other hand is more vulnerable to such artifacts in the training data, since it may have memorized the locked equilibrium configuration corresponding to the input boundary condition. We refer readers to Sec. B.2 for hyperparameters used to train MGN.
4.5 Generalization Experiments
Here we discuss two experiments that assess how well Mass-Spring Net generalizes to novel scenarios not included in training rollouts. Moreover, we demonstrate that despite our neural constitutive model is built from masses and springs, it can be trained to emulate the behavior of source cloths constructed from triangular meshes bearing a different resolution and simulated with FEM principles.
@@ 4.5 Generalization Experiments
Baselines. Tab. 1 compares our method with several recent, open-sourced differentiable cloth or general deformable simulators across key features— if the simulator supports modelling heterogeneous cloths, can simulate surrogate bearing different spatial resolution from the source, if it needs a reference mesh to kick-start training and if it supports simulating long rollouts. We focus exclusively on methods that allow us to model material properties of a stand-alone piece of cloth from kinematic data and optionally, the total mass and physical dimension of the system. While state-of-the-art garment simulators as in (Shao et al., 2025; Li et al., 2024; Grigorev et al., 2023; Santesteban et al., 2022) has achieved stunning results in reconstructing garment kinematics on moving human avatars, we exclude them as they tackle the specific body-and-garment joint fitting problem. In Sec. 4.4 we have made a brief qualitative comparison with Pfaff et al. (2020) in terms of robustness to membrane locking in simulation data. Here we compare, both quantitatively and qualitatively between our surrogate system and the following baseslines, in terms of ability to generalize to new initial and boundary conditions: (i) Neural ODE (Chen et al., 2018) which is an obvious choice for modelling constitutive properties, due to the simplicity and flexibility of its MLP-based architecture — see Fig. 12 in Sec. C for an illustration of the architecture; (ii) GNN (Sanchez-Gonzalez et al., 2020) which can learn to simulate complex, granular materials from point cloud data. We exclude state-of-the-art neural constitutive methods like NCLaw (Ma et al., 2023) from being compared as baselines, since they require premium knowledge such as internal stress in their training curricula to learn complex 3D deformables. Another criterion for our comparisons is the method’s ability to model heterogeneity in cloths, so we have excluded methods like Liang et al. (2019); Li et al. (2022a) which places greater emphasis on modelling homogeneous cloths. We demonstrate that our neural constitutive model is able to generalize better to scenes with novel initial or boundary conditions, and it adapts well to scenes that involve interaction of new objects for long-period simulation that lasts for seconds. We also compare Mass-Spring Net with the two selected baselines in terms of training time and model complexity, measured in number of learnable parameters.
| Time | Mass-Spring Net | GNN | Neural ODE |
@@
Generalization to novel initial and boundary conditions. We train Mass-Spring Net and Neural ODE on rollouts generated from a triangular mesh-based cloth with spatially-varying stiffness: see Fig. 10 (a). The mesh-based cloth is discretized into a particle grid. For running Mass-Spring Net and Neural ODE, we discretize the surrogate cloth into a particle grid, which is relatively prime to the resolution of the source cloth. GNNs are designed to model granular materials, so like Neural ODE it cannot explicitly encode a fixed topological structure, so we do not perform resampling. We then train Mass-Spring Net and the baselines, and use the trained models to reconstruct test-set rollouts with new initial and boundary conditions, visualize the reconstructed rollouts and compute reconstruction error reported in Tab. 2. We report errors from simulating the first , because GNN, like many follow-up work such as Ma et al. (2023) focuses on reconstructing motion of large-scale granular materials over short periods of time, and achieves good performance. Fig. 10 (b) shows the spatial variation in stiffness in the fitted surrogate cloth. Notice that despite the source cloth is built from meshes rather than springs and masses, Mass-Spring Net is still able to recover the two stiffer stripes that run from the top to the bottom, although reconstruction of the top, horizontal stripe and the patch in the middle is sub-optimal. The distribution plots in Fig. 10 (c), (d) shows that the estimated stiffness constants follow a bimodal distribution, which is consistent with the visual in (b). The uniform distribution of damping constant estimates is also expected because triangles in the source cloth have uniform damping, though ideally these estimates should fall in a narrower range. With hyperparamter optimization, Neural ODE was able to learn the initial momentum of the cloth simulation— As Tab. 2 shows, its performance is comparable to GNN and Mass-Spring Net for reconstructing -long rollouts. But for longer rollouts that last for , Neural ODE surrogate completely disintegrates, as shown in Fig. 3. We hypothesize this behavior to be due to lack of topological information available to the model. As the surrogate evolves over time, lacking of topological information allows the particles to roam freely, thereby disintegrating the cloth. Methods that use MLPs to learn dynamics (Ma et al., 2023), do so along with other strong priors as mentioned in the previous section. GNN performs comparably to the mass-spring network in reconstructing the first of rollouts, but its performance declines significantly for -long rollouts. Our method achieves the smallest error in reconstructing long simulations, and we refer readers to Sec. C for detailed analyses of GNN’s performance. In terms of qualitative assessment, we show in Fig. 3 that after simulating for the mass-spring net can attain a steady state closer to the ground-truth than the baselines. We refer readers to Fig. 13 for more qualitative results.
@@
Turning to training time and model complexity: without leveraging domain-specific information, both baselines require more learnable parameters to learn the topology of a cloth: while Mass-Spring Net contains learnable parameters to fit a surrogate system with resolution of , Neural ODE requires over learnable parameters, and GNN contains parameters. We also measured the training time of Mass-Spring Net and the baselines on the same platform, and found that training Mass-Spring Net for two passes requires only minutes. While Neural ODE requires hour of training time which is slightly less than Mass-Spring Net, GNN requires hours which is considerably more. We refer readers to Sec. B.2 for our analysis of GNN’s long training time.
@@
Generalization to novel dynamics. We place the trained surrogate cloth models from the above experiment under a new scene, in which all particles on the surrogate cloth are free to move, and the surrogate cloth starts falling from time until it drops onto another object. We also instantiate the source cloth in the same scene and set it to have the same initial state (position and orientation of the cloth) as the surrogate, and place an identical object below it for the source cloth to drape onto. Figs. 8 and 11 show the steady state after each system settles. Note that as discussed in 4.4, the source cloth used here has the same physical dimension as the surrogate, but bears higher spatial resolution.
@@
Our method produces a significantly better match to the static, draped position of the source cloth with some discrepancy in fine scale details such as fine wrinkles. The Neural ODE model fails to yield a stable result during generalization while GNN methods cannot be easily applied to scenarios with previously unseen objects or contacts in the scene since they roll-out the simulation from an initial set of a few frames. The best way to inject new force-based interactions into these roll outs is still an open question.
5 Conclusion

### v3

# SpringTime: Learning Simulatable Models of Cloth with Spatially-varying Constitutive Properties

@@
Abstract.
Materials used in real clothing exhibit remarkable complexity and spatial variation due to common processes such as stitching, hemming, dyeing, printing, padding, and bonding. Simulating these materials, for instance using finite element methods, is often computationally demanding and slow. Worse, such methods can suffer from numerical artifacts called “membrane locking” that makes cloth appear artificially stiff. Here we propose a general framework, called SpringTime, for learning a simple yet efficient surrogate model that captures the effects of these complex materials using only motion observations. The cloth is discretized into a mass-spring network with unknown material parameters that are learned directly from the motion data, using a novel force-and-impulse loss function. Our approach demonstrates the ability to accurately model spatially varying material properties from a variety of data sources, and immunity to membrane locking which plagues FEM-based simulations. Compared to graph-based networks and neural ODE-based architectures, our method achieves significantly faster training times, higher reconstruction accuracy, and improved generalization to novel dynamic scenarios. Codebase for the paper can be found at https://github.com/ericchen321/springtime.
1. Introduction
@@
In this work, we introduce a learning-based framework, called SpringTime, that addresses all these problems. For run-time simulation, SpringTime uses a simple mass-spring network cloth model that is easy to implement and efficient to simulate in real-time applications. Moreover, we show that SpringTime can effectively learn spatially-varying constitutive properties of cloth from simulated experiments using a more complex and slow model (e.g., an FEM model). Thus SpringTime may be used as a real-time “surrogate” model for a more complex off-line model.
Even though there is a popular perception that mass-spring networks are unsophisticated and less appropriate for modeling continuum mechanics, that is not necessarily the case. As observed by Breen et al. , Hearle and colleagues at Manchester’s renowned Department of Textiles had investigated applications of shell and plate models in the 1980s, and concluded: “But, in dealing with the three-dimensional buckling of textile fabrics, neither the terminology nor the methodology of the established mechanical theory of the bending of plates and shells is of much help because of the limiting assumptions that are made” (Amirbayat and Hearle, 1989). Mass-spring networks could better capture the anisotropic behavior of cloth, and avoid numerical artifacts of FEM discretization of thin shells. We demonstrate, with a series of experiments, that SpringTime is resistant to the membrane locking problem that plagues FEM simulations, while avoiding the complexity of special solutions. Remarkably, we observe that SpringTime behaves better under bending than even the FEM model that it was trained with, thereby achieving superresolution in inference. We illustrate these phenomena and offer some potential explanations.
@@ 2. Related Work
Cloth Simulation and constitutive property estimation. Simulations pertaining to cloth have always been a prime subfield in graphics and simulation research due to their varied industrial applications, from yarn-level simulators by Sperl et al. (2022); Cirio et al. (2014) which specialize in woven cloths, to continuum-level or meshed-based simulators (Shao et al., 2025; Ju et al., 2024; Grigorev et al., 2023; Kairanda et al., 2023; Zhao et al., 2023; Feng et al., 2022; Li et al., 2022b; Santesteban et al., 2022; Bertiche et al., 2022; Li et al., 2022a; Pfaff et al., 2020; Liang et al., 2019; Wang et al., 2011). Many of these works, including Ju et al. (2024); Feng et al. (2022); Wang et al. (2011) focus on accurate constitutive parameter estimates by real-to-sim adaptations using specialized measurement devices; others (Shao et al., 2025; Grigorev et al., 2023; Pfaff et al., 2020) focused more on learning constitutive properties from synthetic data, and demonstrated great generalizability to real-world data (Zheng et al., 2024; Li et al., 2022b, a; Liang et al., 2019), or more specifically, learning garment deformation models jointly with human motion kinematics (Stuyck and Chen, 2023; Zhao et al., 2023; Grigorev et al., 2023; Li et al., 2024; Bertiche et al., 2022; Santesteban et al., 2022). However, cloth or garments chosen as the subject to estimate haven been almost exclusively homogeneous, composed of a single fabric without complex fine-scale patterns and structures; the only exceptions are Stuyck and Chen (2023) and Grigorev et al. (2023), yet both explores fitting specifically of garments on virtual try-on’s. The focus of our work is to address the broader problem of estimating any piece of cloth’s heterogeneous constitutive properties, such as stiffness that varies across space using synthetic data. We distinguish from work like (Stuyck and Chen, 2023; Li et al., 2022a; Pfaff et al., 2020) that our estimation pipeline does not require a reference mesh as input, and from (Zheng et al., 2024; Grigorev et al., 2023; Stuyck and Chen, 2023; Bertiche et al., 2022) that we focus on the broader domain of modelling constitutive properties of a standalone cloth, rather than the specific domain of high-quality reconstruction of garment kinematics on moving human avatars. We show that by fitting a physically correct mass-spring network using our novel loss and training curriculum, we achieve comparable or better accuracy and generalizability than previous approaches but at a significant reduction in training time and model complexity.
@@ 3. Method
We map the predicted spring forces to internal forces on each surrogate particle (Liu et al., 2013),
@@
3.1. Spatial Discretization and Resampling
Given a desired number of surrogate particles , we build our surrogate mass-spring network as follows: we isometrically unroll the source system’s cloth specimen into a rectangle in the plane, then discretize this rectangle with a -vertex regular square grid, as shown in Fig. 5 (left). We place a surrogate spring at each edge of the grid, with the spring rest length determined by the edge length in this rest configuration.
As shown in Fig. 5 (right), the source system landmarks do not necessarily correspond to the surrogate particles (and the resolution of the surrogate particle grid might be much coarser or finer than the density of landmark points). Therefore we must resample the given landmark positions at each frame to positions of the surrogate particles.
@@ 3.1. Spatial Discretization and Resampling
3.2. Neural Constitutive Modeling
@@ 3.3. Simulation and Training
Many alternative loss functions for training the surrogate could also be used, such as position loss (Ma et al., 2023) or acceleration loss (Sanchez-Gonzalez et al., 2020), but in our ablation experiments (Sec. C.2) we found that the force-and-impulse loss works best in our setting.
@@
4. Experiments
4.1. Synthetic Data Generation
Constructing source systems. We build two source systems of cloths modeled under different simulation frameworks: (i) A square-shaped cloth made up of particles and springs. Forces that act upon particles come from springs only; (ii) A square-shaped cloth constructed from a triangular mesh. Internal forces are evaluated based on FEM principles. We use NVIDIA WARP’s FEM-based cloth simulation publicly available on Github (Macklin, 2022b). We discretize each cloth into a particle grid. The choice of varies across experiments; for the experiment in which we compare SpringTime with baselines on generalizability to novel boundary and initial conditions, we use , which is close to the discretization scale picked by recent work such as Shao et al. (2025). We assign different stiffness values to either springs in the mass-spring cloth or triangles in the mesh-based cloth so they have spatially-varying constitutive properties: see Fig. 1.
Generating rollouts with different initial and boundary conditions. To collect ground-truth simulation data, for each cloth we anchor a set of its vertices so they cannot move, and we simulate the deformation of the rest of the cloth over time. The set of vertices being fixed in space, formally defined as anchor neighborhood defines the boundary condition of a rollout; the initial position and velocity of all particles defines its initial condition. For each rollout , we randomly pick two anchor neighborhoods as boundary condition , and we apply a random rotation to the sheet to assign the cloth with initial condition : see Fig. 6. For training SpringTime and the Neural ODE baseline we use rollouts; for testing we use rollouts. Each rollout lasts seconds, during which the cloth falls under the influence of gravity and Rayleigh damping force , where Rayleigh damping constant . We simulate with timestep size of , and this time discretization scheme gives us time steps of data from each rollout.
4.2. Metrics
In some of our experiments, we evaluate SpringTime by applying it to a source system that is itself a simulation of a mass-spring network. In this setting, we can easily compare the surrogate’s estimated spring stiffness and damping parameters to those of the source system. We quantify the error in the reconstructed parameters by computing the root-mean-squared error (RMSE) between their estimated and ground-truth values:
@@ 4.2. Metrics
4.3. Reconstructing Spatially Varying Materials
The goal of this experiment is to show that SpringTime can accurately estimate the stiffness and damping properties of a source cloth with spatially varying stiffness. We use the and metric established in Eq. 13 to assess the quality of our constitutive property estimates, and to use them, we build the source cloth with masses and springs, and we construct a surrogate system with the same resolution as the source cloth. Bending and the complete set of shear springs are included in both the source and the surrogate. This allows us to establish a one-to-one mapping from springs in the source cloth to springs in the surrogate cloth. We also assess the quality of motion reconstruction using the Motion RMSE metric from Eq. 14.
Fig. 7 shows the accuracy of stiffness and damping estimates, and the reconstruction error obtained from a test set which contains -long rollouts. The source cloth we used to generate training and test rollouts is a square-shaped, cloth discretized into a particle grid connected by springs. The source cloth contains springs with three stiffnesses: , and , and all springs share the same damping constant of . Fig. 7 (left) tells that most of our estimated stiffnesses are close to the ground-truth. The damping estimates are also close to the ground-truth on average, despite some variations: see Fig. 7 (middle).
4.4. Robustness against Membrane Locking
@@
To further validate the super-resolution effect of the mass-spring-based surrogate, as shown in Fig. 2 (b), (c), we pin two corners of a square-shaped FEM-simulated mesh along a diagonal, make the cloth fold under the influence of gravity and Rayleigh damping, and sample the cloth’s equilibrium configurations. We compare the equilibrium configurations under two cases: (i) “Aligned”, i.e. folding along the major diagonal that coincides with diagonal edges in the mesh; (ii) “Misaligned”, i.e. folding along the minor diagonal. Under the misaligned case, as we expect, the cloth sags less than the aligned case, and the difference in equilibrium configurations is more apparent on the low-resolution mesh. As shown in Fig. 2 (d), our mass-spring-based surrogate which includes the full set of diagonal and bending springs, and trained on landmark trajectories sampled from the low-resolution mesh in (b), however is robust against this effect, because we can see that at equilibrium configuration it is closer to the high-resolution mesh that suffers less from membrane locking. Further, by comparing Fig. 2 (d), (e) we can see that the surrogate is robust against the pseudo-stiffness caused by membrane locking not by chance, but because it can properly fit to the actual bending resistance of the mesh— the surrogate at equilibrium configuration is closer to the mesh with bending stiffness (d) than the mesh with zero bending stiffness (e).
@@
While both SpringTime and MGN are data-driven methods that encode cloth’s constitutive properties and reconstruct reference motion in inference, what makes SpringTime stands out is its robustness to numerical artifacts that exist in the training data. We believe SpringTime attains the robustness by more accurately estimating the cloth’s constitutive properties, after seeing trajectories sampled in a multitude () rollouts, many of which would not lead to a “locked” equilibrium configuration as in the misaligned case. MGN on the other hand is more vulnerable to such artifacts in the training data, since it tends to memorize the locked equilibrium configuration corresponding to the input boundary condition. We refer readers to Sec. B.2 for hyperparameters used to train MGN.
4.5. Generalization Experiments
Here we discuss two experiments that assess how well SpringTime generalizes to novel scenarios not included in training rollouts. Moreover, we demonstrate that despite our neural constitutive model is built from masses and springs, it can be trained to emulate the behavior of source cloths constructed from triangular meshes bearing a different resolution and simulated with FEM principles.
@@ 4.5. Generalization Experiments
Baselines. Tab. 1 compares our method with several recent, open-sourced differentiable cloth or general deformable simulators across key features— if the simulator supports modelling heterogeneous cloths, can simulate surrogate bearing different spatial resolution from the source, if it needs a reference mesh to kick-start training and if it supports simulating long rollouts. We focus exclusively on methods that allow us to model material properties of a stand-alone piece of cloth from kinematic data and optionally, the total mass and physical dimension of the system. While state-of-the-art garment simulators as in (Shao et al., 2025; Li et al., 2024; Grigorev et al., 2023; Santesteban et al., 2022) has achieved stunning results in reconstructing garment kinematics on moving human avatars, we exclude them as they tackle the specific body-and-garment joint fitting problem. In Sec. 4.4 we have made a brief qualitative comparison with Pfaff et al. (2020) in terms of robustness to membrane locking in simulation data. Here we compare, both quantitatively and qualitatively between our surrogate system and the following baseslines, in terms of ability to generalize to new initial and boundary conditions: (i) Neural ODE (Chen et al., 2018) which is an obvious choice for modelling constitutive properties, due to the simplicity and flexibility of its MLP-based architecture — see Fig. 12 in Sec. C for an illustration of the architecture; (ii) MeshGraphNet (MGN) (Pfaff et al., 2020) which can learn a surrogate model for complex thin-shell materials from point cloud data. We exclude state-of-the-art neural constitutive methods like NCLaw (Ma et al., 2023) from being compared as baselines, since they require premium knowledge such as internal stress in their training curricula to learn complex 3D deformables. Another criterion for our comparisons is the method’s ability to model heterogeneity in cloth, so we have excluded methods like Liang et al. (2019); Li et al. (2022b) which places greater emphasis on modelling homogeneous cloth. We demonstrate that our neural constitutive model is able to generalize better to scenes with novel initial or boundary conditions, and it adapts well to scenes that involve interaction of new objects over extended period of time. We also compare SpringTime with the two selected baselines in terms of training time and model complexity, measured in number of learnable parameters.
| Time | SpringTime | MGN | Neural ODE |
@@
Generalization to novel initial and boundary conditions. We train SpringTime and the baselines on rollouts generated from a triangular mesh-based cloth with spatially-varying stiffness: see Fig. 10 (left). The mesh-based cloth is discretized into a particle grid. For running SpringTime and Neural ODE, we discretize the surrogate cloth into a particle grid, which is relatively prime to the resolution of the source cloth. MeshGraphNet on the other hand requires the source cloth’s default topological structure for training and testing, so we do not perform resampling. We then train SpringTime and the baselines, and use the trained models to reconstruct test-set rollouts with new initial and boundary conditions, visualize the reconstructed rollouts and compute reconstruction error reported in Tab. 2. We see SpringTime outperforms both baselines in terms of motion reconstruction accuracy over long rollouts. Fig. 10 (right) shows the spatial variation in stiffness in the fitted surrogate cloth. Notice that despite the source cloth is built from meshes rather than springs and masses, SpringTime is still able to recover the two stiffer stripes that run from the top to the bottom and the central patch, although reconstruction of the bottom, horizontal stripe is sub-optimal. In terms of qualitative assessment, we show in Fig. 3 that after simulating for SpringTime can attain a steady state closer to the ground-truth than the baselines. We refer readers to Fig. 13 for more qualitative results. We see from the motion RMSE in Tab. 2 and Fig. 3 that Neural ODE surrogate completely disintegrates after of simulation. We hypothesize this behavior to be due to lack of topological information available to the model. As the surrogate evolves over time, lacking of topological information allows the particles to roam freely, thereby disintegrating the cloth. Methods that use MLPs to learn dynamics (Ma et al., 2023), do so along with other strong priors as mentioned in the previous section.
@@
Turning to training time and model complexity: without leveraging domain-specific information, both baselines require more learnable parameters to learn the topology of a cloth: while SpringTime contains learnable parameters to fit a surrogate system with resolution of , Neural ODE requires over learnable parameters, and MGN contains parameters. We also measured the training time of SpringTime and the baselines on the same platform, and found that training SpringTime for two passes requires only minutes, while Neural ODE requires minutes of training time which is slightly less than SpringTime. MGN on the other hand requires considerably more training time: training for iterations requires hours, and we expect training for steps as used in (Pfaff et al., 2020) would require hours.
@@
Generalization to novel dynamics. We place the trained surrogate cloth models from the above experiment under a new scene, in which all particles on the surrogate cloth are free to move, and the surrogate cloth starts falling from time until it drops onto another object. We also instantiate the source cloth in the same scene and set it to have the same initial state (position and orientation of the cloth) as the surrogate, and place an identical object below it for the source cloth to drape onto. Figs. 8 and 11 show the steady state after each system settles. We use Warp (Macklin, 2022b)’s built-in SDF-based collision detector.
@@
SpringTime produces a satisfactory match to the static, draped position of the source cloth, with fine wrinkles that can be difficult to observe if the FEM-simulated source cloth has low resolution. The Neural ODE model fails to yield a stable result during generalization, and methods like MGN cannot be easily applied to scenarios with previously unseen objects or contacts in the scene, since such methods require an input window consisting of multiple frames of kinematic data, and they do not take external forces, contact information or actions as inputs to the model.
5. Conclusion